OpenStudy (anonymous):

If a ball is thrown vertically upward from the roof of 64 foot building with a velocity of 64 ft/sec, its height after t seconds is s(t)=64+64t-16t^2. A) What is the maximum height the ball reaches? (Solved: 128) B) What is the velocity of the ball when it hits the ground (height 0)?

OpenStudy (xapproachesinfinity):

yes what is it that you asking for?

OpenStudy (anonymous):

Sorry, it's my first time doing this and it posted before I wanted it to. I need to find the velocity of the ball when it hits the ground, but I'm not quite sure how to go about doing that.

OpenStudy (xapproachesinfinity):

when the ball hits the ground that mean s(t)=0

OpenStudy (xapproachesinfinity):

he the first part find the vertex like @Zale101 said

OpenStudy (imstuck):

eek...

OpenStudy (imstuck):

put 0 in for s(t) in the position function and solve for t. Then put that t in the velocity equation to find the velocity at height 0.

OpenStudy (xapproachesinfinity):

so you solved the first part let's find the derivative at a point when the s(t)=0 we need to find at what instant t the ball hits the ground

OpenStudy (anonymous):

If I put in 0 for s(t), would that mean plugging it into the quadratic formula? And then plugging that into 64-16t?

OpenStudy (xapproachesinfinity):

Nope solve s(t)=0 solve a quadratic equation

OpenStudy (xapproachesinfinity):

find t=?

OpenStudy (anonymous):

t is 2

OpenStudy (xapproachesinfinity):

so you solve it? can i trust you on that answer?

OpenStudy (anonymous):

Probably not

OpenStudy (xapproachesinfinity):

64+64t-16t^2=0 this is what you solved?

OpenStudy (xapproachesinfinity):

so?

OpenStudy (anonymous):

Do I take the derivative of that and then solve it or do I do 64+64t-16t^2 = 0, change that the 4+4+t^2=0, subtract 8 from both sides, and then take the square root of 8?

OpenStudy (xapproachesinfinity):

nope solve a quadratic equation means find the the zeros for that equation the derivative step is coming later

OpenStudy (xapproachesinfinity):

and by the way it is 4+4t-t^2=0 not what you wrote

OpenStudy (xapproachesinfinity):

solve that

OpenStudy (anonymous):

Okay, for the quadratic formula I go -0.828 and 4.828

OpenStudy (xapproachesinfinity):

did you use this one 4+4t-t^2=0?

OpenStudy (anonymous):

no, i used the first one. should i have used that one?

OpenStudy (xapproachesinfinity):

which one did you call the first one it's the same equatuon 64+64t-16t^2=0 is also 4+4t-t^2=0

OpenStudy (anonymous):

64+64t-16t^2

OpenStudy (xapproachesinfinity):

that's the same as 4+4t-t^2=0

OpenStudy (xapproachesinfinity):

find the two solutions!

OpenStudy (anonymous):

Okay, so I got 10.94 and -6.94

OpenStudy (xapproachesinfinity):

okay! you are positive with that? if so then plug in that into the derivative of s(t)

OpenStudy (xapproachesinfinity):

which is the velocity

OpenStudy (anonymous):

I got -111.04 but it still says I'm wrong? I think I must have missed something.

OpenStudy (xapproachesinfinity):

you didn't do it right www.wolframalpha.com/input/?i=4%2B4x-x%5E2%3D0

OpenStudy (xapproachesinfinity):

you first things was right!

OpenStudy (xapproachesinfinity):

but i was not sure what equation you were using so i had to make sure because you wrote 4+4t+t^2=0 which is not our equation

OpenStudy (anonymous):

I plugged it into the derivative (64-16t)

OpenStudy (xapproachesinfinity):

x=4.8284

OpenStudy (xapproachesinfinity):

this is the value you have to plug in

OpenStudy (xapproachesinfinity):

your derivative is not correct

OpenStudy (xapproachesinfinity):

it should be s'(t)=64-32t

OpenStudy (anonymous):

Oh shoot, sorry I completely missed that. I wrote over that in my notebook. Thanks so much for your help!

OpenStudy (xapproachesinfinity):

welcome^_^

OpenStudy (xapproachesinfinity):

alright i have to go

OpenStudy (xapproachesinfinity):

i think everything is clear now!

OpenStudy (anonymous):

Yes it is! Thanks so much haha. I just completely missed what I wrote as the derivative because I had so much written in my notebook for this problem. Thanks for being patient!

OpenStudy (xapproachesinfinity):

No problem! good luck